I am frustrated and have i tried everthing but can't get it please help me.
The question is :- Solve 7^x = (3^(x+2))*(2^(x-3))
Thank you Darryl

Answer by venugopalramana(3286)

I am frustrated and have i tried everthing but can't get it please help me.
The question is :- Solve 7^x = (3^(x+2))*(2^(x-3))
Thank you Darryl
TAKE LOGS BOTH SIDES
XLOG(7)=(X+2)LOG(3)+(X-3)LOG(2)=XLOG(3)+2LOG(3)+XLOG(2)-3LOG(2)
X{LOG(7)-LOG(3)-LOG(2)}=LOG(3)^2-LOG(2^3)
X{LOG(7/(3*2))=LOG(9/8)
X={LOG(9/8)}/{LOG(7/6)}...I THINK THIS WILL DO ORIF YOU WANT THE ARITHMATIC IT IS
X=5.611